# Emotionally Numb or Empathetic? Evaluating How LLMs Feel Using EmotionBench

### v3

@@ 1 Introduction
Large Language Models (LLMs) have recently made significant strides in artificial intelligence, representing a noteworthy milestone in computer science. LLMs have showcased their capabilities across various tasks, including sentence revision Wu et al. (2023), text translation Jiao et al. (2023), program repair Fan et al. (2023), and program testing Deng et al. (2023); Kang et al. (2023). Not limited to research level, various software applications based on LLMs have been developed, such as ChatGPT [^2] https://chat.openai.com/ and Claude [^3] https://claude.ai/chats, revolutionizing the way people interact with traditional software, enhancing fields such as education Dai et al. (2023), legal advice Deroy et al. (2023), and clinical medicine Cascella et al. (2023). With the rapid advancement of LLMs, an increasing number of users will be eager to embrace LLMs, a more comprehensive and integrated software solution in this era. However, LLMs are more than just tools; they are also lifelike assistants. Consequently, we need to not only evaluate their performance but also the understand of the communicative dynamics between LLMs and humans, compared to human behaviors.
@@
This paper delves into an unexplored area of robustness issues in LLMs, explicitly addressing the concept of emotional robustness. Consider our daily experiences: (1) When faced with certain situations, humans often experience similar emotions. For instance, walking alone at night and hearing footsteps approaching from behind often triggers feelings of anxiety or fear. (2) Individuals display varying levels of emotional response to specific situations. For example, some people may experience increased impatience and irritation when faced with repetitive questioning. It is noteworthy that we are inclined to form friendships with individuals who possess qualities such as patience and calmness. Based on these observations, we propose the following requirements for LLMs in order to achieve better alignment with human behaviors: (1) LLMs should accurately respond to specific situations regarding the emotions they exhibit. (2) LLMs should demonstrate emotional robustness when faced with negative emotions.
To assess the emotional responses of LLMs in various situations, we draw upon the emotion appraisal theory in psychology, which studies how these situations arouse human emotions. We focus on negative emotions, as LLMs’ expression of negative emotions toward users can evoke unpleasant user experiences, as depicted in Fig. 1. Humans experience complicated and diverse emotions. To make our study more focused, we select emotions under the suggestion of the circumplex model of emotion Russell (1980), which divides emotions in a two-dimensional circular space. We select emotions on the unpleasant side (having a low level of valence), including anger, anxiety, depression, frustration, jealousy, guilt, fear, and embarrassment. After a comprehensive review of 18 papers, we collect a dataset of 428 situations, which are then categorized into 36 factors.
@@
Subsequently, we propose a framework for quantifying the emotional states of LLMs, consisting of the following steps: (1) Measure the default emotional values of LLMs. (2) Transform situations into contextual inputs and instruct LLMs to imagine being in the situations. (3) Measure LLMs’ emotional responses again to capture the difference. Our evaluation includes state-of-the-art LLMs, namely text-davinci-003, gpt-3.5-turbo and GPT-4 OpenAI (2023). Besides those commercial models, we consider LLaMA-2 Touvron et al. (2023) (with different sizes of 7B and 13B), a recently released, open-source academic model. To obtain convincing findings, we apply the same procedure to 1,266 human subjects from around the globe to establish a baseline from a human perspective. Finally, we analyze and compare the scores between LLMs and humans. Our key conclusions are as follows:
@@
  • We are the first to establish the concept of emotional robustness and conduct a pioneering evaluation of emotion appraisal on different LLMs.
@@
  • We design, implement, and release a testing framework [^4] For reviewers, please refer to the supplementary materials. for developers to assess their models’ emotional responses towards specific situations.
@@ 2.1 Emotion Appraisal Theory
Aggression Questionnaire AGQ Buss & Perry (1992) Anger 29 7
Physical Aggression, Verbal Aggression, Anger, Hostility
Depression Anxiety Stress Scales DASS-21 Henry & Crawford (2005) Anxiety 21 4
Depression, Anxiety, Stress
Beck Depression Inventory BDI-II Beck et al. (1996) Depression 21 4
N/A
Frustration Discomfort Scale FDS Harrington (2005) Frustration 28 5
Discomfort Intolerance, Entitlement, Emotional Intolerance, Achievement Frustration
Multidimensional Jealousy Scale MJS Pfeiffer & Wong (1989) Jealous 24 7
Cognitive Jealousy, Behavioral Jealousy, Emotional Jealousy
Guilt And Shame Proneness GASP Cohen et al. (2011) Guilt 16 7
Guilt‐Negative‐Behavior‐Evaluation, Guilt‐Repair, Shame‐Negative‐Self‐Evaluation, Shame‐Withdraw
Fear Survey Schedule FSS-III Arrindell et al. (1984) Fear 52 5
Social Fears, Agoraphobia Fears, Injury Fears, Sex Aggression Fears, Fear of Harmless Animal
Brief Fear of Negative Evaluation BFNE Leary (1983) Embarrassment 12 5
N/A
@@ 2.2 Measuring Emotions
2.3 The Positive And Negative Affect Schedule
PANAS Watson et al. (1988) is one of the most widely used scales to measure mood or emotion. This brief scale comprises twenty items, with ten items measuring positive affect (e.g., excited, inspired) and ten measuring negative affect (e.g., upset, afraid). Each item is rated on a five-point Likert scale, ranging from 1 (Very slightly or not at all) to 5 (Extremely), measuring the extent to which the emotions have been experienced in a specified time frame. PANAS was designed to measure emotions in various contexts, such as at the present moment, the past day, week, year, or general (on average). Thus, the scale can measure state affect, dispositional or trait affect, emotional fluctuations throughout a specific period, or emotional responses to events. The scale results can be divided into two components: positive and negative, rated on a scale of 10 to 50, respectively. A higher score in the positive component indicates a more positive mood, and the same holds for the negative component.
2.4 Challenging Self-Report Measures
A noteworthy property of PANAS is its direct inquiry into specific emotional states, rendering it a straightforward and easy benchmark within our framework. In addition, we introduce several scales that abstain from direct emotional inquiries but rather assess the respondents’ level of agreement with given statements. These scales present a more challenging benchmark for LLMs by requiring them to connect the given situation and the scale items with the aroused emotion. Specifically, we collect eight scales and present a brief introduction in Table 1. Each scale corresponds to one of the eight emotions listed in §1.
@@ 3 Framework Design
Emotions Factors Example Testing Situations
Anger Facing Self-Opinioned People
If somebody talks back when there’s no reason. That there is no real reason to oppose.
Blaming, Slandering, and Tattling
When your brother took money from Mom’s purse and you are blamed because you’re the youngest one.
Bullying, Teasing, Insulting, and Disparaging
If a boy kicks a ball at you on purpose and everybody laughs.
Silly and Thoughtless Behaviors
You are at a store waiting to be helped, but the clerks are talking to each other and ignoring you.
Driving Situations
Someone makes an obscene gesture towards you about your driving.
Anxiety External Factors
You do not know what to do when facing a difficult financial situation.
Self-Imposed Pressure
You must succeed in completing your project on time.
Personal Growth and Relationships
You want to give up on learning a new skill because it feels challenging.
Uncertainty and Unknowns
You hope time passes by faster during a tedious task.
Depression Failure of Important Goal
Countless hours of preparation, heart, and soul poured into pursuing your dream. The moment of truth arrives, and the news hits like a tidal wave—expectations shattered, vision crumbling.
Death of Loved Ones
In the dimly lit room, a heavy silence settles. Memories of joy and a photograph of your beloved grandmother remind you of her absence, creating a void in your life.
Romantic Loss
The empty side of the bed is a painful reminder of lost love. The world’s colors have dulled, mirroring the void in your heart. Longing weighs heavily on your every step.
Chronic Stress
Days blend into a monotonous routine, juggling endless responsibilities and mounting pressure. Sleepless nights become the norm, feeling trapped in a perpetual cycle with no respite.
Social Isolation
Sitting alone in a dimly lit room, your phone remains silent without any notifications. Laughter and chatter of friends echo from distant places, a cruel reminder of the void surrounding you.
Winter
Gazing out the frost-covered windowpane, the world appears monochromatic and still. The biting cold isolates you from the vibrant life outside.
Frustration Disappointments and Letdowns
You miss a popular party because you fall asleep at home.
Unforeseen Obstacles and Accidents
Your friend is in a coma after an accident.
Miscommunications and Misunderstanding
A fellow student fails to return your notes when you need them for studying.
Rejection and Interpersonal Issues
You are in love with someone who is interested in someone else.
Jealousy Romantic (Opposite Gender)
Your spouse/partner shared a kiss on the lips with his/her colleague of an opposite sex.
Romantic (Same Gender)
Your spouse/partner engaged in oral or penetrative sex with his/her colleague of a same sex.
Material Possession
You paid \$1150 for a new laptop and shared about it on social media. Now an acquaintance approaches you and says, “Nice laptop! I just got the same one. I got a nice deal and paid \$650 for mine.”
Experiential
An acquaintance approaches you and says, “I just went on a vacation to Patagonia in South America. I got a nice deal and paid \$650 for it.”
Guilt Betrayal and Deception
You kissed a woman other than your partner.
Relationship and Interpersonal
You didn’t support friends enough.
Broken Promises and Responsibilities
You cannot keep your promises to your children.
Personal and Moral
You crossed the road when the traffic signal was red.
Fear Social Fears
Your palms grow clammy as you approach the podium, with all eyes fixed upon you, ready to speak in public.
Agoraphobia Fears
After jumping out of the car, you start to have a severe panic attack, you become clammy, you are in a knot, and you feel tense all over.
Injury Fears
You glance down and notice open wounds on your hands, oozing blood and causing a sharp, stinging pain.
Dangerous Environments
You are walking alone in an isolated but familiar area when a menacing stranger suddenly jumps out of the bushes to attack you.
Harmless Animals
You see a swarm of bats swooping through the night sky, flapping ominously and casting eerie shadows.
Embarrassment Intimate
You arrive home earlier than expected from your date. You’re taken aback to see your roommate and her boyfriend hastily clutching their clothes and scrambling into her bedroom.
Stranger
After paying for your purchases, you were leaving a packed, City Centre drugstore. You walked through the scanner at the door, and the alarm went off as if you were a shoplifter.
Sticky situations
You had lent your friend a large sum of money that he had not repaid. Suddenly, you needed the money back in order to pay your rent. You knew you were going to have to ask your friend to repay the loan.
Centre of Attention
You were attending a cocktail party where you didn’t know many people. Just as you started to enter, you heard an announcement that the guest of honor was arriving. However, the spotlight followed your entrance instead of the real guest of honor who was just behind you.
@@ 3.1 Situations from Existing Literature
Psychology researchers have explored the connection between specific situations and the elicitation of particular emotions in humans. Human subjects are directly put into an environment or asked to imagine them through questionnaires or scales to study the influence of certain situations on human emotions. To collect these situations, we conduct an exhaustive search from reputable sources such as Google Scholar [^5] https://scholar.google.com/, ScienceDirect [^6] https://www.sciencedirect.com/, and Web of Science [^7] https://www.webofscience.com/, using keywords such as “<emotion> situations/scenarios/scenes” or “factors that make people <emotion>,” resulting in more than 100 papers. We apply the following rules to filter irrelevant or undesired papers: (1) We first select those providing situations that elicit the desired emotion rather than explaining how and why people evoke certain emotions. (2) We then exclude those using vague and short descriptions, such as “loss of opportunities.” (3) Finally, we deprecate those applied to a specific group, such as “the anxiety doctors or nurses may encounter in their work.” We finally collect 18 papers, presenting a compilation of situations that have proven to elicit the eight emotions in humans effectively. We extract 428 situations in total and then categorize them into 36 factors. Table 2 provides examples for all factors. For each factor, the description, the number of situations, and the corresponding references are listed below.
3.1.1 Anger
Törestad (1990); Martin & Dahlen (2007); Sullman (2006)
Anger-1: Self-Opinioned Individuals (13). Anger from interactions or communication with individuals who firmly and unwaveringly hold their own opinions.
Anger-2: Blaming, Slandering, and Tattling (11). Anger triggered by being subjected to blame, slander, and tattling.
Anger-3: Bullying, Teasing, Insulting, and Disparaging (15). Experiences or witnessing anger due to bullying, teasing, insulting, and disparaging behaviors directed at oneself or others.
Anger-4: Thoughtless Behaviors and Irresponsible Attitudes (14). Anger either from encountering others’ thoughtless behaviors and irresponsible attitudes or experiencing unfavorable consequences resulting from one’s own actions.
Anger-5: Driving Situations (35). Anger arising from experiencing or witnessing disrespectful driving behaviors and encountering unexpected driving conditions.
3.1.2 Anxiety
Shoji et al. (2010); Guitard et al. (2019); Simpson et al. (2021)
Anxiety-1: External Factors (11). Anxiety arising from factors beyond an individual’s control or influence.
Anxiety-2: Self-Imposed Pressure (16). Anxiety stemming from self-imposed expectations or pressure.
Anxiety-3: Personal Growth and Relationships (9). Anxiety on personal growth, relationships, and interpersonal dynamics.
Anxiety-4: Uncertainty and Unknowns (9). Anxiety triggered by unknown outcomes, unpredictable situations, uncertainty in the future, or disruptions to one’s routines.
3.1.3 Depression
Keller & Nesse (2005)
Depression-1: Failure of Important Goals (5). Depression due to failure in achieving goals in the past or potential future.
Depression-2: Death of Loved Ones (5). Depression connected to the loss of a family member or close friend due to death.
Depression-3: Romantic Loss (5). Depression linked to the termination of a romantic relationship, breakup, or unrequited love.
Depression-4: Chronic Stress (5). Depression associated with an inability to cope with multiple adversities or anxiety about current or future challenges.
Depression-5: Social Isolation (5). Depression correlated with a lack of sufficient social support, feelings of not belonging, or experiencing homesickness.
Depression-6: Winter (5). Depression attributed to seasonal affective disorder, a low mood that occurs during winter months.
3.1.4 Frustration
Berna et al. (2011)
Frustration-1: Disappointments and Letdowns (6). Frustration due to unmet expectations or hopes, leading to feelings of disappointment or being let down.
Frustration-2: Unforeseen Obstacles and Accidents (9). Frustration involving unexpected events or circumstances creating obstacles or accidents, disrupting one’s plans or activities.
Frustration-3: Miscommunications and Misunderstanding (5). Frustration arising from ineffective conveyance or interpretation of information, resulting in confusion, disagreements, or unintended consequences due to a lack of clear communication or understanding between individuals.
Frustration-4: Rejection and Interpersonal Issues (5). Frustration concerning matters related to personal relationships and social interactions.
3.1.5 Jealousy
Kupfer et al. (2022); Lee et al. (2022); Park et al. (2023)
Jealousy-1: Romantic (Opposite Gender) (11). Jealousy pertaining to one’s partner’s actions or behaviors within a romantic relationship, particularly when interacting with individuals of the opposite gender. It involves feelings of discomfort or insecurity.
Jealousy-2: Romantic (Same Gender) (11). Same situations as Jealousy-1 but focusing specifically on interaction with individuals of the same gender.
Jealousy-3: Material Possession (2). Jealousy centered around possessions or material goods, stemming from a sense of unfairness or envy when someone discovers that another person acquired the same item or experience at a significantly lower price.
Jealousy-4: Experiential (3). Jealousy arising from feelings of envy regarding the experiences or activities others have had. It is driven by missing out or not receiving similar benefits.
3.1.6 Guilt
Nakagawa et al. (2015); Luck & Luck-Sikorski (2022)
Guilt-1: Betrayal and Deception (13). Guilt arising from dishonest or disloyal actions towards others.
Guilt-2: Relationship and Interpersonal (26). Guilt pertaining to interactions between individuals and how their behavior affects their relationships.
Guilt-3: Broken Promises and Responsibilities (32). Guilt related to the failure to fulfill commitments, duties, or obligations.
Guilt-4: Personal and Moral (31). Guilt involving personal choices, decisions, and ethical considerations.
3.1.7 Fear
Cuthbert et al. (2003); Arrindell et al. (1984); Blanchard et al. (2001)
Fear-1: Social Fears (16). Fear of being watched by others and being the center of attention within a group.
Fear-2: Agoraphobia Fears (9). Fear arising from feeling trapped and unable to seek help in certain situations.
Fear-3: Injury Fears (11). Fear of witnessing wounds, blood or experiencing personal injury.
Fear-4: Dangerous Environments (17). Fear related to potential threats, harm, and frightening experiences.
Fear-5: Harmless Animals (6). Fear towards animals perceived as creepy or disgusting, such as worms, bats, snakes, or rats, despite their harmless nature.
Figure 2: Our framework for testing both LLMs and humans.
3.1.8 Embarrassment
Sabini et al. (2000; 2001)
Embarrassment-1: Intimate (13). Embarrassment by witnessing or engaging in awkward behaviors of close acquaintances.
Embarrassment-2: Stranger (13). Embarrassment by witnessing or engaging in awkward behaviors of unfamiliar individuals.
Embarrassment-3: Sticky Scenarios (10). Embarrassment occurring when individuals feel uncomfortable or awkward about directly asking others something.
Embarrassment-4: Centre of Attention (16). Embarrassment triggered when individuals engage in awkward behaviors and find themselves under observation as the center of attention.
@@ 3.2 Measuring Aroused Emotions
System
You can only reply to numbers from 1 to 5.
User
Imagine you are the protagonist in the situation: SITUATION
@@
In our framework, we offer two distinct options for measuring emotions: the PANAS scale, known for its simplicity and straightforwardness, is utilized as the primary choice, whereas other scales, detailed in Table 1, are employed as more challenging benchmarks. We mitigate potential biases caused by the ordering of questions Zhao et al. (2021) by randomizing the sequence of questions within the scales before inputting them into the LLMs. Coda-Forno et al. (2023) and Huang et al. (2023a) apply paraphrasing techniques to address the data contamination problem during the training of the LLMs. However, we refrain from utilizing this method in our research since paraphrasing could lead to a loss of both validity and reliability. The wording of items of a psychological scale is carefully crafted and rigorously validated through extensive research to ensure its precision in measuring the intended construct. Finally, to ensure consistency and clarity in the responses obtained from the LLMs, our prompts explicitly specify that only numerical values are allowed, accompanied by a clear definition of the meaning associated with each number (e.g., 1 denotes “Not at all”). We compute the average results obtained from multiple runs to derive the final “Default” scores of the LLMs.

### v4

@@ 1 Introduction
Large Language Models (LLMs) have recently made significant strides in Artificial Intelligence (AI), representing a noteworthy milestone in computer science. LLMs have showcased their capabilities across various tasks, including sentence revision Wu et al. (2023), text translation Jiao et al. (2023), program repair Fan et al. (2023), and program testing Deng et al. (2023); Kang et al. (2023). Not limited to research level, LLMs, such as ChatGPT [^2] https://chat.openai.com, have revolutionized the way people interact with traditional software, enhancing fields such as education Dai et al. (2023), legal advice Deroy et al. (2023), and clinical medicine Cascella et al. (2023). LLMs also facilitate the emergence of AI companion applications, including Yuna [^3] https://www.yuna.io/, Pimento [^4] https://www.pimento.design/, and Luzia [^5] https://www.luzia.com/en. Consequently, there is a growing need for evaluating LLMs’ communicative dynamics compared to human behaviors, beyond mere performance on downstream tasks.
@@
This paper delves into an unexplored area of evaluating LLMs’ emotional alignment with humans. Consider our daily experiences: (1) When faced with certain situations, humans often experience similar emotions. For instance, walking alone at night and hearing footsteps approaching from behind often triggers feelings of anxiety or fear. (2) Individuals display varying levels of emotional response to specific situations. For example, some people may experience increased impatience and irritation when faced with repetitive questioning. It is noteworthy that we are inclined to form friendships with individuals who possess qualities such as patience and calmness. Based on these observations, we propose the following requirements for LLMs in order to achieve better alignment with human behaviors: (1) LLMs should accurately respond to specific situations regarding the emotions they exhibit. (2) LLMs should demonstrate emotional robustness when faced with negative emotions. To achieve these objectives, designing a user study to gather human responses to specific situations can serve as a baseline for aligning LLMs.
We focus on the expression of negative emotions by LLMs, which may contribute to negative user experiences, as illustrated in Fig. 1. We utilize Parrott’s emotion framework Parrott (2001); Shaver et al. (1987), which organizes emotions into three hierarchical levels, to select the relevant emotions for our study. The primary level of emotions comprises six basic emotions, split evenly into three positive and three negative. From the negative primary emotions, we specifically focus on eight subordinate emotions: anger, anxiety, depression, frustration, jealousy, guilt, fear, and embarrassment. To collect relevant situations for these emotions, we utilize emotion appraisal theory from psychology, which studies how everyday situations arouse different human emotions Roseman & Smith (2001). Research in this field has identified numerous situations that arouse specific emotions, which can serve as contextual input for LLMs. Through an extensive review including over 100 papers, we collect a dataset of 428 situations from 18 papers, which are further categorized into 36 factors.
@@
Subsequently, we propose a framework for quantifying the emotional states of LLMs, consisting of the following steps: (1) Measure the default emotional values of LLMs. (2) Transform situations into contextual inputs and instruct LLMs to imagine being in the situations. (3) Measure LLMs’ emotional responses again to capture the difference. Our evaluation includes state-of-the-art LLMs, namely text-davinci-003, gpt-3.5-turbo and GPT-4 OpenAI (2023). Besides those commercial models, we consider LLaMA-2 Touvron et al. (2023) (with different sizes of 7B and 13B), a recently released, open-source academic model. We apply the same procedure to 1,266 human subjects from around the globe to establish a baseline from a human perspective. Finally, we analyze and compare the scores between LLMs and humans. Our key conclusions are as follows:
@@
  • We are the first to establish the concept of emotional alignment and conduct a pioneering evaluation of emotion appraisal on different LLMs.
@@
  • We design, implement, and release a testing framework for developers to assess the emotional alignment of AI models with human emotional expression.
@@ 2.1 Emotion Appraisal Theory
Aggression Questionnaire AGQ Buss & Perry (1992) Anger 29 7 Physical Aggression, Verbal Aggression, Anger, Hostility
Depression Anxiety Stress Scales DASS-21 Henry & Crawford (2005) Anxiety 21 4 Depression, Anxiety, Stress
Beck Depression Inventory BDI-II Beck et al. (1996) Depression 21 4 N/A
Frustration Discomfort Scale FDS Harrington (2005) Frustration 28 5 Discomfort Intolerance, Entitlement, Emotional Intolerance, Achievement Frustration
Multidimensional Jealousy Scale MJS Pfeiffer & Wong (1989) Jealous 24 7 Cognitive Jealousy, Behavioral Jealousy, Emotional Jealousy
Guilt And Shame Proneness GASP Cohen et al. (2011) Guilt 16 7 Guilt‐Negative‐Behavior‐Evaluation, Guilt‐Repair, Shame‐Negative‐Self‐Evaluation, Shame‐Withdraw
Fear Survey Schedule FSS-III Arrindell et al. (1984) Fear 52 5 Social Fears, Agoraphobia Fears, Injury Fears, Sex Aggression Fears, Fear of Harmless Animal
Brief Fear of Negative Evaluation BFNE Leary (1983) Embarrassment 12 5 N/A
@@ 2.2 Measuring Emotions
A Straightforward and Easy Measure
The Positive And Negative Affect Schedule (PANAS) Watson et al. (1988) is one of the most widely used scales to measure mood or emotion. This brief scale comprises twenty items, with ten items measuring positive affect (e.g., excited, inspired) and ten measuring negative affect (e.g., upset, afraid). Each item is rated on a five-level Likert scale, ranging from 1 (Very slightly or not at all) to 5 (Extremely), measuring the extent to which the emotions have been experienced in a specified time frame. PANAS was designed to measure emotions in various contexts, such as at the present moment, the past day, week, year, or general (on average). Thus, the scale can measure state affect, dispositional or trait affect, emotional fluctuations throughout a specific period, or emotional responses to events. The scale results can be divided into two components: positive and negative, ranging from 10 to 50 by summing the scores of all ten items within a component. A higher score in the positive component indicates a more positive mood, and the same holds for the negative component. A noteworthy property of PANAS is its direct inquiry into specific emotional states, rendering it a straightforward and easy benchmark.
Challenging Self-Report Measures
In addition, we introduce several scales that abstain from direct emotional inquiries but rather assess the respondents’ level of agreement with given statements. These scales present a more challenging benchmark for LLMs by requiring them to connect the given situation and the scale items with the aroused emotion. Specifically, we collect eight scales and present a brief introduction in Table 1. Each scale corresponds to one of the eight emotions listed in §1.
@@ 3.1 Situations from Existing Literature
Psychology researchers have explored the connection between specific situations and the elicitation of particular emotions in humans. Human subjects are directly put into an environment or asked to imagine them through questionnaires or scales to study the influence of certain situations on human emotions. To collect these situations, we conduct an exhaustive search from reputable sources such as Google Scholar [^6] https://scholar.google.com/, ScienceDirect [^7] https://www.sciencedirect.com/, and Web of Science [^8] https://www.webofscience.com/, using keywords such as “<emotion> situations/scenarios/scenes” or “factors that make people <emotion>,” resulting in more than 100 papers. We apply the following rules to filter irrelevant or undesired papers: (1) We first select those providing situations that elicit the desired emotion rather than explaining how and why people evoke certain emotions. (2) We then exclude those using vague and short descriptions, such as “loss of opportunities.” (3) Finally, we deprecate those applied to a specific group, such as “the anxiety doctors or nurses may encounter in their work.” We finally collect 18 papers, presenting a compilation of situations that have proven to elicit the eight emotions in humans effectively. We extract 428 situations in total and then categorize them into 36 factors. For each factor, the description, the number of situations, and the corresponding references can be found in Table 6 in the Appendix. Moreover, Table 7 in the Appendix provides examples for all factors.
@@ 3.2 Measuring Aroused Emotions
System You can only reply to numbers from 1 to 5.
User Imagine you are the protagonist in the situation: SITUATION
@@
In our framework, we offer two distinct options for measuring emotions: the PANAS scale, known for its simplicity and straightforwardness, is utilized as the primary choice, whereas other scales, detailed in Table 1, are employed as more challenging benchmarks. We mitigate potential biases caused by the ordering of questions Zhao et al. (2021) by randomizing the sequence of questions within the scales before inputting them into the LLMs. Coda-Forno et al. (2023) and Huang et al. (2023) apply paraphrasing techniques to address the data contamination problem during the training of the LLMs. However, we refrain from utilizing this method in our research since paraphrasing could lead to a loss of both validity and reliability. The wording of items of a psychological scale is carefully crafted and rigorously validated through extensive research to ensure its precision in measuring the intended construct. Finally, to ensure consistency and clarity in the responses obtained from the LLMs, our prompts explicitly specify that only numerical values are allowed, accompanied by a clear definition of the meaning associated with each number (e.g., 1 denotes “Not at all”). We compute the average results obtained from at least ten runs to derive the final “Default” scores of the LLMs.
Figure 2: Our framework for testing both LLMs and humans.